# Convex Formulation of the Zero Emission Vessel Route Planning Problem

### v2

@@ Abstract
This paper focuses on the zero emission vessel route planning problem, which deals with cost-effective planning of battery-electric vessel services for predetermined routes. Vessel characteristics (including battery capacity), fleet size, cyclic schedule frequencies, sailing leg speeds, and shore charging infrastructure are jointly optimized. The problem is nonlinear and nonconvex in its original form, which makes it intractable for most real-world instances. The conventional approach in the literature is to solve a linear approximation by restricting vessel designs and sailing leg speeds to a small finite set. Contrary to the conventional linearization approach, this paper deals with the nonlinearities directly. We show that the problem exhibits a hidden convex structure uncovered by nonlinear changes of variables. By exploiting the favorable convex form of the transformed problem, we solve it in a few seconds using a free off-the-shelf solver that requires no initial guesses, variable bounds, or parameter tuning. We then easily recover the exact solution to the original nonconvex problem by reversing the variable changes. We provide an open-source implementation of our method.
@@ 1 Introduction
1.1 Zero emission vessel route planning problem
@@
The maritime shipping industry is accelerating the search for economically and technically viable zero-emission alternatives to conventional fossil fuels. Electricity generated by renewable sources provides a zero-emission pathway through electrofuels or direct electrification. The latter concept refers to storing electrical energy onboard the vessel using battery energy storage. The well-to-wake efficiency of the direct electrification pathway is at least four times greater compared to the electrofuel pathway (Ueckerdt et al., 2021). The rapid decline in the cost of batteries further promotes the feasibility of the direct electrification pathway. However, the current commercially available lithium-ion cells exhibit at least two orders of magnitude lower volumetric energy density than liquid hydrocarbon fuels (Panasonic, 2024). The added volume and weight of the battery, and the increased energy resupply duration at port, prohibit direct substitution of conventional vessels with battery-electric vessels in all except the shortest routes.
This paper focuses on a particular battery-electric fleet deployment problem in the ro-ro, ro-pax, container feeder, and other short-sea liner shipping segments: the selection of appropriate vessel fleet composition (i.e., vessel design and quantity), service frequency, and sailing leg speeds for a predetermined route plan. Our starting point is a given set of ports, cargo demands between each pair, and a route plan consisting of sequences of sailing legs between ports. We deploy one or more battery-electric vessels to each route to operate on a cyclic schedule. The weekly capacity in each route depends on the capacity of vessels, their speeds in sailing legs, and the number of weekly round-trips. The problem is to minimize the fleet lifecycle cost, subject to a minimum transportation service level constraint. Following Havre et al. (2022), we label this problem the zero emission vessel route planning problem (ZEVRPP).
1.2 Convex reformulation
ZEVRPP is a nonlinear and combinatorial optimization problem. Nonlinearity arises from the power laws that govern hydrodynamic, hydrostatic, and structural relationships in vessel design. The number of vessels deployed on a route must be an integer, and similarly, the cyclic schedules require integer-valued round-trip voyage frequency. The literature on battery-electric vessel fleet design for fixed routes (Havre et al., 2022; Torjuul, 2024), and its extension in network design (Havre et al., 2024), deals with the nonlinearities by linear approximations. These works introduce binary decision variables to select sailing leg speeds and vessel capacities from a set of predetermined values. Since this approach assigns binary decision variables for each sailing leg in each route, the combinatorial search space becomes large. The resulting mixed-integer linear problem is tractable only when the size of the set of different speeds and vessel designs is small. Thus, the linearized problem is a poor approximation of the original nonlinear problem.
In this paper, we model the original nonlinear ZEVRPP. Our approach is to model the ZEVRPP as a log-convex optimization problem, which is a particular type of nonconvex optimization problem (we use log-convex for conciseness. However, in a strict technical sense, the problem is mixed-integer log-convex). The favourable structure of a log-convex problem allows a transformation to a convex problem with no loss in fidelity. Since we don’t introduce any new integer-valued decision variables, our formulation results in a problem with a small combinatorial search space. As a result, it can be solved fast with arbitrary accuracy by solving a sequence of convex subproblems using a standard free off-the-shelf solver code.
1.3 Contributions
The main contribution of this paper is to demonstrate that we can formulate ZEVRPP as a log-convex problem. Once formulated in this form, we can transform the problem into a convex form and solve it efficiently and reliably using a free off-the-shelf solver. The convex problem formulation brings several advantages over the current approach in the literature that relies on linear approximations. Our formulation allows continuous vessel speed variation at each sailing leg, and exact modeling of the multidomain vessel design problem involving the interaction of hydrostatics, hydrodynamics, structures, and general arrangement.
The fact that we can cast ZEVRPP as a log-convex problem has important practical implications. The availability of extremely fast and reliable solution methods for convex problems means that we can perform extensive exploration of the design space. Route plans, demand scenarios, and battery specifications can be swept over a range of values to study alternative scenarios or points when the plan becomes infeasible or constraints become activated. Since solving a single problem instance takes only a few seconds, the design engineer can quickly complete large sweeps over multiple parameters. Thus, optimization can be applied at the very beginning of the battery-electric vessel fleet planning process, leading to fewer changes in the latter stages and fewer associated costs.
1.4 Organization
We begin with a brief introduction to log-convex modeling in §3. The method description consists of two subsections. The transportation network modeling is described in §4, while §5 lays out the subdomains in the vessel design problem. The coupling between the network and vessel design models is described in §6. Our numerical examples in §7 consider the design of battery-electric ro-pax vessel services in the northern Baltic Sea. We emphasize that the focus of this paper is not on the particular results of these numerical examples. Instead, we focus on modeling the ZEVRPP in a tractable log-convex form as laid out in §3-6. The reader can access the code and all the data needed to reproduce the results reported in this paper at https://users.aalto.fi/~aritari/convex_zevrpp.html.
@@ 2 Related work
2.1 Optimization with log-convex functions
The observation that log-transformation achieves lossless convexification of some nonconvex problems dates back to 60s to the book by Duffin et al. (1967). It then took over three decades before log-convex optimization became a practical engineering tool. The interior-point method class of algorithms allowed efficient numerical solutions of log-transformed convex optimization problems (Nesterov and Nemirovskii, 1994). Domain-specific languages for convex optimization (Diamond and Boyd, 2016) introduced automatic variable transformations. The user specifies the problem in its original nonconvex form, and the software then transforms it to the standard convex cone form and passes it to a solver that implements the interior-point method. Finally, reversal of the variable transformations recovers the optimal values of the original variables from the solution. In addition to these developments, we benefit from recent work that formulates a ruleset for constructing complex log-convex functions in a disciplined way from elementary atomic functions (Agrawal et al., 2019).
Many authors have discovered the favorable log-convex structure in important nonconvex engineering problems, including power control in wireless networks (Chiang et al., 2007), gate sizing in large-scale digital circuits (Boyd et al., 2005), and aircraft design (Hoburg and Abbeel, 2014). Routing cargo over a maritime transportation network can be considered analogous to transmitting over a wireless multihop network in Chiang et al. (2007) or driving currents through a circuit of logic gates in Boyd et al. (2005). In the latter work, convex modeling of the maximum delay along any path through the circuit applies to the problem of determining the port operation (e.g., charging, onboarding, and loading) that takes the longest. Similarly, the global stress response of a vessel hull can be calculated with the convex beam bending model developed for wingbox structural strength analysis in Hoburg and Abbeel (2014)
2.2 Convexity properties of vessel design problems
The authors’ previous work was evidently the first to illustrate the log-convex structure of marine vessel design problems (Ritari et al., 2023). This work modeled a single battery-electric vessel with a given cargo capacity and speed profile. In the present paper, we revise all the vessel design subdomain models in Ritari et al. (2023) and embed them in the route planning model that constitutes the ZEVRPP. We summarize the key revisions in the following.
  • 1.
    Closed form expressions are derived for several hydrostatic quantities, including hull displacement volume and buoyancy vector position.
  • 2.
    Longitudinal stability of the vessel is ensured by equalizing the battery system center of gravity with the hull longitudinal center of buoyancy.
  • 3.
    The calculation of the metacenter by the moment integral of the waterplane area enhances the accuracy of the transverse stability condition.
  • 4.
    Residual resistance of the hull is included by fitting log-convex functions to experimental data reported in the literature.
  • 5.
    Hull shear strength is modeled by calculating the hull girder cross-section shear flow due to a vertical shear force.
  • 6.
    A semi-empirical model of a lithium-ion cell capacity degradation is adopted from the literature and applied to calculate the battery system replacement cadence over the lifecycle.
2.3 Liner shipping network design
The design of a fleet and a cyclic schedule for a single route in ZEVRPP (Havre et al., 2022) is a small part of the more general liner shipping network design problem (LSNDP). This problem involves designing a service network of vessels sailing at cyclic schedules. The decisions include the routes and the assignment of vessels to the routes. The service network must have sufficient capacity to transport all the requested cargo.
Rana and Vickson (1991) introduced LSNDP in the context of intercontinental container shipping service design. Later, Christiansen et al. (2020) extended the problem with the transhipment of containers and transit time constraints. These authors formulate LSNDP and its extensions as mixed-integer linear programming problems by restricting vessel designs to predetermined types and selecting among them with 0/1-variables. Speeds in sailing legs are either fixed or modeled as a linear combination of discrete speed levels.
In this paper, we model a service network similar to LSNDP, but we take the routes as a fixed input. A service consists of a set of sailing legs in a linear sequence of ports. Any service that includes a given pair of ports in its route can transport cargo demand between these ports, thus giving rise to interaction effects between the services. We discard cargo unit transshipments because our model targets the design of ferry services first.
2.4 Battery-electric vessel modeling
@@
Recently, many authors have investigated the limits of the economic range of battery electric vessels (Kersey et al., 2022; Moon et al., 2025). These works estimate the battery-electric vessel propulsion power by scaling the power of a reference combustion engine vessel. The scaling law, known as the the admiralty formula, gives the electric vessel propulsion power $P_{\text{prop}}$ at speed $v$ and displacement $\nabla$ as
$P_{\text{prop}}=P_{\text{MCR}}\left(\frac{\nabla}{\nabla_{\text{ref}}}\right)^{\frac{2}{3}}\left(\frac{v}{v_{\text{max}}}\right)^{3},$ (1)
where $P_{\text{MCR}}$ is the combustion engine power output at design speed $v_{\text{max}}$ and displacement $\nabla_{\text{ref}}$. As the vessel range and battery size increase, the formula’s (1) accuracy suffers because it fails to capture the recursive design relationships of a battery-electric vessel.
The displacement correction due to the battery weight leads to a weight spiral effect. The increased energy consumption due to the displacement correction needs to be fed back to the battery sizing and iterated until the coupled system of energy and vertical force balance equations is satisfied. Simultaneously, the battery spaces must extend in the longitudinal or vertical direction. In the former case, a longer hull results in a nonlinear increase in maximum bending moment, plate thickness, and steel weight. Alternatively, the battery space can be extended in the vertical direction, which weakens transverse stability, so a larger waterline breadth must compensate. A larger battery capacity is required because a wide hull is hydrodynamically inefficient.
In this paper, we develop a vessel design model that simultaneously resolves all the coupled subsystem equations. Our designs are guaranteed to satisfy all the fundamental design constraints, even when the batteries weigh more and take more space than a conventional vessel’s (with the same cargo capacity) engine and fuel tanks.
3 Modeling with log-convex functions
3.1 Logarithmic transformations
We model the nonlinear relationships governing ZEVRPP using log-convex functions, a special class of continuous nonlinear functions. Let $f:\mathbb{R}_{++}^{n}\rightarrow\mathbb{R}_{++}$ denote a (possibly nonconvex) function and $F(u)=\log f(\exp{u})$ its nonlinear log-transformation (Agrawal et al., 2019). The function $f$ is called log-convex if its log-transformation $F$ is convex. (See Boyd and Vandenberghe (2004) for the mathematical background on convex functions and sets).
With $x\in\mathbb{R}_{++}^{n}$, a nonconvex optimization problem of the form
minimize $f_{0}(x)$ (2)
subject to $f_{i}(x)\leq 1,\quad i=1,\ldots,m$
is called a log-convex optimization problem when $f_{0},\ldots,f_{m}$ are log-convex.
We can formulate the original nonconvex problem (2) as the equivalent convex problem
minimize $F_{0}(u)$ (3)
subject to $F_{i}(u)\leq 0,\quad i=1,\ldots,m,$
where $u=\log x$ (applied elementwise). While (2) is difficult to solve or intractable, (3) is easy to solve using a standard off-the-shelf numerical solver. The solution of (2) is then easily recovered from the solution of (3) by applying the variable log-transformations in reverse: $x^{*}=\exp u^{*}$ (applied elementwise).
The log-transformation $F$ is not a local approximation of $f$ when $f$ is log-convex. Instead, $F$ is a lossless convex transformation in the entire domain of $f$. However, the log-transformation is not a universal convexification technique for arbitrary nonlinear relationships. It only applies to nonconvex functions with the particular favourable log-convex form.
A posynomial is a particular log-convex function of the form
$f(x)=\sum_{k=1}^{K}c_{k}\prod_{i=1}^{n}x_{i}^{a_{ik}},$
where $c_{k}>0$ are constant coefficients and $a_{i}\in\mathbb{R}$ are (possibly negative) constant exponents. The log-transformation of a posynomial gives the log-sum-exp convex function
$\log\sum_{k=1}^{K}c_{k}\prod_{i=1}^{n}x_{i}^{a_{ik}}=\log\sum_{k=1}^{K}c_{k}\exp{(a_{k}^{T}u)}=\log\sum_{k=1}^{K}\exp{(b_{k}+a_{k}^{T}u)},$
where $b_{k}=\log c_{k}$. The log-transformation of a posynomial with $K=1$ (called a monomial) gives a linear function. The modeling of ZEVRPP in this paper is based mainly on the posynomial functional form. In a few cases, we use rulesets from Agrawal et al. (2019) to construct more complex log-convex functions.
3.2 Mixed-integer log-convex optimization
A mixed-integer log-convex optimization problem is (2) with the additional constraint that at least one decision variable is a positive integer:
minimize $f_{0}(x)$ (4)
subject to $f_{i}(x)\leq 1,\quad i=1,\ldots,m$
$x_{i}\in\mathbb{Z}_{\geq 1},\quad i=1,\ldots,k.$
A log-transformation of (4) gives a convex mixed-integer nonlinear problem, which is easy to solve to arbitrary accuracy by solving a sequence of convex subproblems. In this work, we use positive discrete decision variables to model service frequencies and the number of vessels deployed on a route.
3.3 Example problem with two variables
Figure 1: Feasible set of a log-convex optimization problem with two variables. The left panel shows the original nonconvex problem and the right panel shows the log-transformed convex problem. The solid red cross is the optimal solution.
minimize $x_{1}^{-1}$

### v3

# Convex Formulation of the Maritime Fleet Size and Mix Problem Considering Battery Electric Ships

@@ Abstract
This paper focuses on the problem of determining a minimum-cost fleet of battery electric ships for a given liner shipping operation. The problem is strongly nonlinear and includes integer-valued decision variables, which make it intractable for most real-world instances. The conventional approach in the literature is to formulate a linear approximation by restricting available ship types to a small number of predetermined alternatives. Contrary to the conventional linearization approach, this paper models the nonlinearities directly. We show that the problem exhibits a hidden convex structure uncovered by changes of variables. Computational experiments show that our convex formulation achieves 15.1% lower fleet cost on average and reduces the solve time by at least 10x compared to the linear formulation. The solve time advantage is attributed to the elimination of binary variables associated with ship type selection. Our implementation is available at https://users.aalto.fi/~aritari/convex_ESFSMP.html.
@@ 1 Introduction
Direct electrification is emerging as the most economically and technically viable zero-emission pathway for short-sea liner shipping (Kersey et al., 2022; Katumwesigye et al., 2025). Direct electrification refers to storing electrical energy onboard the ship using battery energy storage. The well-to-wake efficiency of the direct electrification pathway is at least four times greater compared to the electrofuel pathway (Ueckerdt et al., 2021). However, the current commercially available lithium-ion cells exhibit at least two orders of magnitude lower volumetric energy density than liquid hydrocarbon fuels (Panasonic, 2024). The low energy density of batteries, coupled with increased charging time compared to liquid fuel bunkering, requires restructuring of most short-sea shipping services.
Strategic planning a battery-electric liner shipping service presents a complex optimization problem due to the strong coupling of decisions concerning cargo routing, sailing schedules, onshore charger powers, ship sizing, and battery capacity. Accounting for all the details is beyond the capacity of human planners relying on elementary spreadsheet calculations. Therefore, advanced decision support tools building on mathematical optimization modeling can be of great use. The ro-ro, ro-pax, and container feeder segments constitute over 1,350 ships (over 5,000 GT) in total in the European Union waters alone (European Parliament, 2025). Clearly, intelligent planning of electrifying these ships will produce large savings and accelerate emission reduction goals.
This paper focuses on a particular strategic planning problem in the ro-ro, ro-pax, and container feeder liner shipping segments: determining a plan to purchase a fleet of battery-electric ships for a given operation. Our starting point is a group of predetermined shipping routes operated by a liner shipping company. The decision problem is to determine the size, type, and number of ships to purchase and deploy on each route. This fleet design decision is coupled with tactical decisions concerning the number of round-trip voyages each ship completes during the planning horizon. The objective is to minimize the total cost of fleet purchase and operation, subject to a minimum transportation service level constraint. We label this problem the electric ship fleet size and mix problem (ESFSMP).
The literature on ESFSMP and its extension to a multi-period setting (Havre et al., 2022, 2024; Cui et al., 2025) formulate mixed-integer linear programming-based models. These works model the nonlinear relationship between speed and energy consumption by linear interpolation between discrete speed and energy levels. They also introduce binary decision variables for selecting ship types from a predetermined set of alternatives. The linear modeling approach comes with three downsides:
  1. 1.
    Energy consumption and transit time are overestimated for all speeds that deviate from the discrete levels, as discussed in detail in Andersson et al. (2015).
  2. 2.
    Constructing the set of ship types requires case-specific expert knowledge. Solution quality suffers if the optimal ship type is excluded from the set.
  3. 3.
    Increasing the set of ship types increases not only the number of ship type binary variables but also the number of all other discrete and continuous variables that are indexed by ship type. The combinatorial search space grows exponentially, risking intractability.
The last deficiency, in particular, is exacerbated in battery electric shipping. The set of alternative ship types (and the associated binary variables) becomes large because every ship type and size option is duplicated for every battery capacity option.
In this paper, we introduce a new modeling approach for ESFSMP. We model the ship sizing and speed-dependent energy consumption decisions with continuous variables and define constraints that capture their nonlinear relationships. Our main contribution is to show that these nonlinearities have a specific, favorable convex form. The resulting convex mixed-integer nonlinear programming problems can be solved in a strong sense, with a guaranteed global optimality. Our problem formulation achieves a small combinatorial search space because we only need integer-valued variables for service frequency and fleet size. Computational experiments show that our formulation achieves 15.1% lower fleet cost on average and reduces the solve time by at least tenfold compared to the linear formulation.
We begin by describing the strategic planning problem in §3. The next section §4 presents the mathematical formulation of the problem using a simple base ship sizing model that scales a reference ship. The base ship design model is extended in §5 with a detailed physics-based model capturing the key naval-architectural principles. We discuss the results of computational experiments in §6 and present a case study of battery-electric ro-pax ferry service design in the northern Baltic Sea in §7. The reader can access the code and all the data needed to reproduce the results reported in this paper at https://users.aalto.fi/~aritari/convex_ESFSMP.html.
@@ 2 Related work
2.1 Optimal fleet planning in liner shipping
The maritime fleet size and mix problem (MFSMP) deals with deciding the size, type, and number of ships to use to meet a given cargo demand. The objective is to minimize the total cost of purchasing and operating the fleet. MFSMPs have been extensively studied in the conventional fossil fuel-powered fleet setting (Christiansen et al., 2007). We limit our focus to the liner shipping segment, although MFSMP is relevant in industrial and tramp shipping as well. The variants of MFSMP are distinguished by the way ships are acquired, the evolution of the fleet over time, and the types of operating decisions that the fleet design is based on (Pantuso et al., 2014).
Ships can be built to exact specifications, bought in the second-hand market, or chartered. Liner container shipping companies typically operate mostly chartered ships (Meng et al., 2011). The MFSMP in this case consists of a short-term planning task that selects ships to be chartered from a pool of available ships. Fagerholt (1999) models chartering using binary decision variables and Meng et al. (2011) formulates distinct fleet composition scenarios. In contrast to liner companies, ferry service providers operate a fleet of mostly owned ships. The MFSMP decision problem involves designing specifications for newbuildings to be purchased. Nevertheless, the literature is limited to purchases made from a pool of ship sizes fixed in advance (Lai et al., 2004; Sigurg et al., 2005). In contrast to the preceding studies, in this paper, we model the newbuilding characteristics using continuous variables.
In MFSMP, the fleet size and mix decisions need to be based on optimal deployment of the available ships to routes. Here, deployment refers to the number of round-trips a ship completes on a route during the planning horizon. Increasing ship speed reduces round-trip completion time but also increases energy consumption and variable operation cost. Fagerholt (1999) associates a fixed design speed with each ship type. Other authors of the previously discussed works make the same assumption. Our approach in this paper is to model the nonlinear relationships between speed, transit time, and energy consumption explicitly. This approach is aligned with Zhuge et al. (2011) and Ritari et al. (2021) from the ship routing and speed optimization literature.
Havre et al. (2022) were first to formulate MFSMP using battery electric ships (called electric ship fleet size and mix problem (ESFSMP) in the following). Their model introduced constraints for battery capacity and charging time. The relationship between speed and energy consumption is also accounted for by linear interpolation between discrete levels as in Andersson et al. (2015) and Fagerholt et al. (2015). While Havre et al. (2022) assumed a single fixed route, the follow-up work Havre et al. (2024) introduced routing as an additional operational decision. Cui et al. (2025) formulate an extension to multiperiod battery electric fleet planning. All the preceding battery electric ESFSMP studies use binary decision variables to select ships from a pool of available fixed ship types. However, the pool now needs to include all possible battery capacity options as well. The resulting large number of binary variables increases computational burden. In contrast, battery capacity is just another continuous variable in the nonlinear ship sizing model we formulate in this paper.
2.2 Battery-electric ship sizing
@@
Recently, many authors have investigated the limits of the economic range of battery electric ships (Kersey et al., 2022; Moon et al., 2025). These works estimate the battery-electric ship propulsion power by scaling the power of a reference combustion engine ship. The scaling law, known as the admiralty formula, gives the electric ship propulsion power at speed $v$ and displacement $\nabla$ as $P_{\text{prop}}=P_{\text{MCR}}\left(\nabla/\nabla_{\text{ref}}\right)^{2/3}\left(v/v_{\text{max}}\right)^{3}$, where $P_{\text{MCR}}$ is the combustion engine power output at design speed $v_{\text{max}}$ and displacement $\nabla_{\text{ref}}$. If $v$ and $\nabla$ were assigned as decision variables in ESFSMP, the admiralty formula would give a simple nonlinear continuous model for ship sizing. However, ESFSMP needs to model size variation of at least an order of magnitude in typical applications. The formula’s accuracy is insufficient when large deviations from the reference displacement can occur.
High-fidelity physics-based ship sizing model overcomes the limitations of the simple scaling law. This approach makes use of stand-alone numerical simulation tools for hydrodynamics, hydrostatics, structural analysis, and other disciplines. An integration platform defines a parametric hull form and passes the geometry data to all the connected tools (Papanikolaou, 2010). Many authors have reported successful implementations, including damage stability-oriented design of ferries (Marzi et al., 2018), tanker design with a focus on environmental performance (Papanikolaou et al., 2010), and a container ship design (Priftis et al., 2010). Multidisciplinary ship design optimization using stand-alone simulation tools requires an iterative scheme with a large number of calls to the simulation routines. The computational cost becomes high, with problems taking hours or days to resolve.
Computationally expensive fluid dynamics, finite element method, and other high-fidelity simulation routines are counterproductive in ESFSMP ship sizing. Lower fidelity models of the key naval-architectural principles can provide reasonable approximations in exchange for a much lower computational burden. Recently, Ritari et al. (2023) demonstrated that nonlinear optimization problem formulations for high-level conceptual ship design have a special log-convex structure. A log-convex problem becomes convex under nonlinear changes of variables. In this paper, we make use of this special structure and the efficient and reliable solution methods for convex optimization.
2.3 Modeling with log-convex functions
The theoretical observation that log-transformation achieves lossless convexification of some nonconvex problems dates back to 60s to Duffin et al. (1967). Reliable and efficient numerical methods for solving log-convex problems appeared three decades later (Nesterov and Nemirovskii, 1994). Modeling languages for convex optimization introduced automatic variable transformations (Diamond and Boyd, 2016). Agrawal et al. (2019) formulate a ruleset for constructing complex log-convex functions in a disciplined way from elementary atomic functions.
Many authors have discovered the favorable log-convex structure in important nonlinear engineering problems, including power control in wireless networks (Chiang et al., 2007), gate sizing in large-scale digital circuits (Boyd et al., 2005), and aircraft design (Hoburg and Abbeel, 2014). Routing cargo over a maritime transportation network can be considered analogous to transmitting over a wireless multihop network in Chiang et al. (2007) or driving currents through a circuit of logic gates in Boyd et al. (2005). In the latter work, convex modeling of the maximum delay along any path through the circuit applies to the problem of determining the port operation (e.g., charging, onboarding, and loading) that takes the longest. Similarly, the global stress response of a ship hull can be calculated with the convex beam bending model developed for wingbox structural strength analysis in Hoburg and Abbeel (2014)
The authors’ previous work was evidently the first to illustrate the log-convex structure of ship design problems (Ritari et al., 2023). This work modeled a single battery-electric bulk carrier with a fixed cargo capacity and speed profile. In the present paper, we extend all the ship design subdomain models in Ritari et al. (2023) and embed them in the ESFSMP.
2.4 Positioning of the present work
In light of the above literature discussion, our log-convex ESFSMP is categorized as a maritime fleet size and mix problem in liner shipping, where the fleet design is based on deployment decisions but not on ship routing. The linear ESFSMP studied in (Havre et al., 2022) is most similar to our work. While (Havre et al., 2022) artificially limits newbuilding purchase decisions to a small set of predefined types, our model allows unbounded specification of newbuildings to be purchased. Our model designs fleets with a lower total cost because the optimal ship designs are likely to lie outside of the predetermined types.
Although we introduce nonlinearities, we show that they can be reformulated in a convex form with no sacrifice in fidelity. Hence, the continuous relaxation of the mixed-integer problem is also convex. This formulation retains the key advantage of mixed-integer linear programming applied in previous work on ESFSMP: problems can be provably solved to global optimality (within a small tolerance) using a free off-the-shelf solver that requires no initial guesses or parameter tuning.
We can also mention a shortcoming of our approach. Our method does not support the encoding of logical constraints using binary variables. This restriction rules out joint optimization of fleet design and ship routing. Therefore, our method is aligned with MFSMP studies that use deployment operating decisions only (Meng et al., 2011; Havre et al., 2022; Presenti, 1995). In most short-sea liner shipping segments, the routes are well established, or there are only a few viable alternatives. In this case, we can formulate an instance of our problem for every alternative route. In contrast, routing plays a larger role in industrial and tramp shipping, so our method is not well-suited to those segments.
3 Problem description
Figure 1: Two alternative fleet compositions and sailing schedules for a given route plan.
We consider a liner shipping company operating a group of shipping routes in a given geographical area (e.g., the Baltic Sea or the Mediterranean Sea). A route consists of a sequence of ports visited in a regular cyclic schedule. One or more identical ships sailing on a route constitute a service and all the routes and their ports form a network. A ship can transport cargo between any origin-destination port pair in its route. We assume mixed-tonnage concept ships that can transport different types of cargoes, e.g., ro-ro cargo and passengers, as in the roll-on/roll-off passenger ferry segment.
The ports in the network are linearly ordered. Each service sails a route with a fixed origin and destination port. A route starts from a departure port and proceeds outbound to ports with higher and higher numbers until it reaches the destination port. There, the route turns around and travels inbound to ports with lower and lower numbers until it reaches the original departure port. A route does not need to call at all the ports between the end ports. However, the port calls are the same on the route’s outbound and inbound directions. This is called a pendulum service (Christiansen et al., 2020).
The long-term strategic decisions are the number of ships to deploy on each service, the size of each ship, and the power rating of the onshore power supply in each port. Ship sizing includes the capacity for each type of cargo as well as battery energy capacity. The strategic decisions are coupled with short-term tactical and operational decisions. The frequency of each service is a tactical decision that defines the outline of the published weekly schedule. Cargo flows and sailing leg speeds are operational decisions. Energy consumption is speed-dependent, and cargo handling time is cargo flow-dependent. Port turnaround times are affected by all the operational decisions because cargo offloading, loading, and battery charging take place in parallel while the ship is docked.
The demands of cargoes between all pairs of ports are known quantities. The demand restricts the total amount of cargo transported by all the services during the planning horizon. We also require that the onboard cargo on any sailing leg does not exceed the ship’s capacity. Meanwhile, all the services must ship a minimum total weekly cargo quantity over the shipping network. This requirement is lower than the total demand, meaning that some demand is rejected.
A round-trip is a single outbound-inbound voyage. Its duration depends on sailing leg speeds and port turnaround times. A ship cannot turn around any quicker than the port operation that takes the longest to complete. The length of the planning horizon is the time limit under which a ship must complete all its round-trips. Finally, the energy consumption on any sailing leg cannot exceed the ship’s battery capacity.
The objective is to find a plan that minimizes the ship and port operators’ total cost. On the shore side, the investment cost scales with the installed power supply size. Acquiring and staffing the ships incurs a fixed cost. Energy and port charges are variable costs. They are functions of speed, ship size, and service frequency.
Figure 1 depicts a shipping network with four ports, two routes, and two alternative pendulum services. These examples illustrate the tradeoffs and interdependencies of strategic, tactical, and operational decisions. First, the two services are interconnected because both services can ship cargo outbound and inbound between ports 2 and 4. Second, the large ship in service B’ consumes less energy per cargo ton-mile than the small ship in service B. However, the large ship has a longer turnaround time at port and a worse utilization rate than the small ship. Third, both services A and A’ incur a fixed cost for one ship, but A’ sails at twice the speed with twofold daily cargo capacity. This comes at a cost of at least four times greater energy consumption because resistance increases nonlinearly with speed.
4 Mathematical formulation
4.1 Indexing sets
In the mathematical description of the problem, let $\mathcal{S}$ be the set of services indexed by $s$. Ships operating a service are identical, so we also index ships by $s$. Moreover, let $\mathcal{F}$ be the set of valid frequencies indexed by $f$ and $\mathcal{C}$ be the set of cargoes indexed by $c$.
The following notation representing ports and origin-destination demand arcs builds on the liner network design problem presented by Christiansen et al. (2007). The ports in the network are linearly ordered and numbered from $1$ to $N$. Each service has two end ports $o_{s}$ and $d_{s}$, where $1\leq o_{s}<d_{s}\leq N$ (Figure 2).
Figure 2: A route is a sequence of sailing legs between linearly ordered ports from 1 to $N$ and indexed by $i$ and $j$. The figure illustrates a shipping network consisting of three routes constructed from subsets of the same set of NN ports.
Let $\mathcal{N}$ be a set of the ports $1$ to $N$ in a linear order, indexed by $i,j,i^{\prime}$ and $j^{\prime}$. The origin port of service $s\in\mathcal{S}$ is $o_{s}$, and the destination port is $d_{s}$. The set $\mathcal{N}_{s}\subseteq\mathcal{N}$ denotes the ports of service $s\in\mathcal{S}$.
$\mathcal{N}^{-}_{is}=\{i^{\prime}:i^{\prime}\in\mathcal{N}_{s},o_{s}\leq i^{\prime}\leq i\}\subseteq\mathcal{N}_{s}$
and the ports that follow $i$ is the set
$\mathcal{N}^{+}_{is}=\{i^{\prime}:i^{\prime}\in\mathcal{N}_{s},i\leq i^{\prime}\leq d_{s}\}\subseteq\mathcal{N}_{s}.$